# The Unbounded Denominators Conjecture

### v1

@@ 7.2. Unbounded denominators for the solutions of certain ODEs
Let $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ be an integer coefficients formal power series solution of $L(f)=0$, where $L$ is a linear differential operator without singularities on $\mathbf{P}^{1}\smallsetminus\{0,1/16,\infty\}$. If the $x=0$ local monodromy of $L$ is finite, then $f(x)$ is algebraic, and more precisely, the function $f(\lambda(\tau)/16)$ on $\mathbf{H}$ is automorphic for some congruence subgroup $\Gamma(N)$ of $\mathrm{SL}_{2}(\mathbf{Z})$.
@@
The condition in Theorem 7.2.1 that the linear differential operator $L$ has a finite local monodromy at $x=0$ is essential for algebraicity. The canonical and explicit transcendental example, which is given in [And04, Appendix, A.5] and we have already mentioned in our introduction § 1.1, is the Gauss hypergeometric series
@@
a weight one modular form for the congruence group $\Gamma_{0}(4)$. The modularity streak is not an accident: more generally, we may reversely start with any congruence modular form of a weight $k>0$, such as for instance Ramanujan’s (discriminant) weight $12$ modular form $\Delta(\tau)=q\prod_{n=1}^{\infty}(1-q^{n})^{24}\in q\mathbf{Z}\llbracket q\rrbracket$, and express it formally into a power series in $x:=\lambda(\tau)/16$, using $\mathbf{Z}\llbracket q\rrbracket=\mathbf{Z}\llbracket x\rrbracket$ as in §1.1. It is then a classical fact, cf. Stiller [Sti84] or Zagier [Zag08, § 5.4], that the resulting formal power series fulfills a linear ODE on $\mathbf{P}^{1}\smallsetminus\{0,1/16,\infty\}\cong Y(2)$ of order $k+1$ and monodromy group $\mathrm{Sym}^{k}\,\mathrm{SL}_{2}(\mathbf{Z})\hookrightarrow\mathrm{SL}_{k+1}(\mathbf{Z})$.
It remains to us an open question whether a complete description of all integral solutions $f\in\mathbf{Z}\llbracket x\rrbracket$ on dropping the $x=0$ finite local monodromy condition in Theorem 7.2.1 should arise in this way from a classical congruence modular form expressed into a holonomic function in $x=\lambda/16$. We formulate the precise statement in Question 7.4.1 below.
@@ 7.3. Vector-valued modular forms
The transition is as follows. By taking the componentwise product $F(\tau)^{12}(\lambda(\tau)/16\Delta(\tau/2))^{k}$ and restricting the full modular group to its subgroup $\Gamma(2)$, we reduce to the case $k=0$ of local systems on $Y(2)\cong\mathbf{P}^{1}\smallsetminus\{0,1/16,\infty\}$. Without loss of generality upon passing to a factor, we may assume that local system to be irreducible. The holomorphic vector bundle with integrable connection admitting $F$ for its horizontal sections is indeed meromorphic at the cusps of $Y(2)$ due to the existence of the $q$-expansion of $F$. Hence $F$ is a solution of a rank-$n$ system of first-order linear homogeneous ODEs over $\mathbf{Q}[\lambda,1/\lambda,1/(1-\lambda)]$. By the theorem of the cyclic vector, see [DGS94, § III.4], there is an irreducible linear differential operator $L$ over $\mathbf{Q}(\lambda)$ without singularities on $Y(2)=\mathrm{Spec}\,\mathbf{Z}[\lambda,1/\lambda,1/(1-\lambda)]$ and such that all $n$ component functions $F_{1},\ldots,F_{n}$ are formal solutions of the linear homogeneous ODE $L(f)=0$. Since one of these (namely, $F_{j}$) has a $\lambda=0$ formal expansion in $\mathbf{Z}\llbracket q\rrbracket=\mathbf{Z}\llbracket\lambda/16\rrbracket$, Chudnovsky’s theorem [DGS94, Theorem VIII.1.5] implies that $L$ satisfies the Galočkin (finite global operator height $\sigma(L)<\infty$) condition [DGS94, VII.2.(2.3) on page 227], hence by the Bombieri–André theorem [DGS94, Theorem VII.2.1], $L$ satisfies the Bombieri (finite generic global inverse radius $\rho(L)<\infty$) condition [DGS94, VII.2.(2.1) on page 226], and is therefore globally nilpotent. At this point Katz’s local monodromy theorem [Kat70] (see also [DGS94, Theorem III 2.3 (ii)]) proves that $L$ has quasi-unipotent local monodromies. Now by the semisimplicity condition on $\rho\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$ in Definition 7.3.1, it follows that in fact the $x=0$ local monodromy of $L$ has finite order. The result then follows on applying Theorem 7.2.1 to $f(x)=F_{j}(\tau)$. ∎
@@ 7.4. Some questions and concluding remarks
If one drops the semisimplicity stipulation on $\rho\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$ in the definition of a vector-valued modular form, the resulting structure has been named a logarithmic vector-valued modular form by Knopp and Mason [KM11]. They also do arise in conformal field theories, termed logarithmic (in place of rational). See, for example, Fuchs–Schweigert [FS19]. But now by Remark 7.2.3 the components of a weight zero vector-valued modular form with bounded denominators can certainly be transcendental over $\mathbf{C}(\lambda)$. Still the examples there are classical (congruence) modular forms, except of a higher weight. The following is an extension of the unbounded denominators problem over to the logarithmic setting. It remains outside the scope of our method as far as we could see.
Question 7.4.1.
@@
If a component $F_{j}(\tau)$ of a logarithmic vector-valued modular form for $\mathrm{SL}_{2}(\mathbf{Z})$ has a $\mathbf{Z}\llbracket q\rrbracket$ Fourier expansion, is $F_{j}(\tau)$ a classical modular form for a congruence subgroup?

### v2

@@ 7.2. Unbounded denominators for the solutions of certain ODEs
Let $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ be an integer coefficients formal power series solution of $L(f)=0$, where $L$ is a linear differential operator without singularities [^3] Similarly to 2.0.4, the proof allows for singularities on $\mathbf{P}^{1}\smallsetminus\{0,1/16,\infty\}$ provided their local monodromy is trivial. on $\mathbf{P}^{1}\smallsetminus\{0,1/16,\infty\}$. If the $x=0$ local monodromy of $L$ is finite, then $f(x)$ is algebraic, and more precisely, the function $f(\lambda(\tau)/16)$ on $\mathbf{H}$ is automorphic for some congruence subgroup $\Gamma(N)$ of $\mathrm{SL}_{2}(\mathbf{Z})$.
@@
The condition in Theorem 7.2.1 that the linear differential operator $L$ has a finite local monodromy at $x=0$ is essential for algebraicity. The canonical and explicit transcendental example, which is given in [And04, Appendix, A.5] and we have already mentioned in our introduction § 1.1, is the Gauss hypergeometric series or complete elliptic integral of the first kind
@@
a weight one modular form for the congruence group $\Gamma(2)$. The modularity streak is not an accident: more generally, to get $\mathbf{Z}\llbracket x\rrbracket$ holonomic functions on $\mathbf{P}^{1}\smallsetminus\{0,1/16,\infty\}$ with infinite $x=0$ local monodromy, we may reversely start with any congruence modular form of a weight $k>0$, such as for instance Ramanujan’s (discriminant) weight $12$ modular form $\Delta(\tau)=q\prod_{n=1}^{\infty}(1-q^{n})^{24}\in q\mathbf{Z}\llbracket q\rrbracket$, and express it formally into a power series in $x:=\lambda(\tau)/16$, using $\mathbf{Z}\llbracket q\rrbracket=\mathbf{Z}\llbracket x\rrbracket$ as in §1.1. It is then a classical fact, cf. Stiller [Sti84] or Zagier [Zag08, § 5.4], that the resulting formal power series fulfills a linear ODE on a finite étale cover of $\mathbf{P}^{1}\smallsetminus\{0,1/16,\infty\}\cong Y(2)$, of order $k+1$ and monodromy group commensurable with $\mathrm{Sym}^{k}\,\mathrm{SL}_{2}(\mathbf{Z})\hookrightarrow\mathrm{SL}_{k+1}(\mathbf{Z})$.
It remains to us an open question whether a complete description of all integral solutions $f\in\mathbf{Z}\llbracket x\rrbracket$ on dropping the $x=0$ finite local monodromy condition in Theorem 7.2.1 should arise in this way from a classical congruence modular form expressed into a holonomic function in $x=\lambda/16$. We formulate the precise statement in Question 7.4.3 below.
@@ 7.3. Vector-valued modular forms
The transition is as follows. By taking the componentwise product $F(\tau)g(\tau)(\lambda(\tau)/16\Delta(\tau/2))^{\frac{k+k^{\prime}}{12}}$, where we choose a non-zero scalar-valued modular form $g(\tau)\in\mathbf{Z}[[q]]$ of weight $k^{\prime}$ such that $12\mid k+k^{\prime}$, and restricting the full modular group to its subgroup $\Gamma(2)$, we reduce to the case $k=0$ of local systems on $Y(2)\cong\mathbf{P}^{1}\smallsetminus\{0,1/16,\infty\}$. Without loss of generality upon passing to a factor, we may assume that local system to be irreducible. The holomorphic vector bundle with integrable connection admitting $F$ for its horizontal sections is indeed meromorphic at the cusps of $Y(2)$ due to the existence of the $q$-expansion of $F$. Hence $F$ is a solution of a rank-$n$ system of first-order linear homogeneous ODEs over $\mathbf{Q}[\lambda,1/\lambda,1/(1-\lambda)]$. By the theorem of the cyclic vector, see [DGS94, § III.4], there is an irreducible linear differential operator $L$ over $\mathbf{Q}(\lambda)$ without singularities on $Y(2)=\mathrm{Spec}\,\mathbf{Z}[\lambda,1/\lambda,1/(1-\lambda)]$ and such that all $n$ component functions $F_{1},\ldots,F_{n}$ are formal solutions of the linear homogeneous ODE $L(f)=0$. Since one of these (namely, $F_{j}$) has a $\lambda=0$ formal expansion in $\mathbf{Z}\llbracket q\rrbracket=\mathbf{Z}\llbracket\lambda/16\rrbracket$, Chudnovsky’s theorem [DGS94, Theorem VIII.1.5] implies that $L$ satisfies the Galočkin (finite global operator height $\sigma(L)<\infty$) condition [DGS94, VII.2.(2.3) on page 227], hence by the Bombieri–André theorem [DGS94, Theorem VII.2.1], $L$ satisfies the Bombieri (finite generic global inverse radius $\rho(L)<\infty$) condition [DGS94, VII.2.(2.1) on page 226], and is therefore globally nilpotent. At this point Katz’s local monodromy theorem [Kat70] (see also [DGS94, Theorem III 2.3 (ii)]) proves that $L$ has quasi-unipotent local monodromies. Now by the semisimplicity condition on $\rho\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$ in Definition 7.3.1, it follows that in fact the $x=0$ local monodromy of $L$ has finite order. The result then follows on applying Theorem 7.2.1 to $f(x)=F_{j}(\tau)$. ∎
@@ 7.4. Some questions and concluding remarks
If one drops the semisimplicity stipulation on $\rho\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$ in the definition of a vector-valued modular form, the resulting structure has been named a logarithmic vector-valued modular form by Knopp and Mason [KM11]. They also do arise in conformal field theories, termed logarithmic (in place of rational). See, for example, Fuchs–Schweigert [FS19]. But now by Remark 7.2.3 the components of a weight zero vector-valued modular form with bounded denominators can certainly be transcendental over $\mathbf{C}(\lambda)$. Still the examples there are classical (congruence) modular forms, except of a higher weight. In Question 7.4.3 below we give an extension of the unbounded denominators problem over to the logarithmic setting. It remains outside the scope of our method as far as we could see. Before stating this question, we recall some basic facts concerning quasi-modular forms. Recall that the ring of quasi-modular forms $\widetilde{M}(\Gamma)$ for $\Gamma\subset\mathrm{SL}_{2}(\mathbf{Z})$ may be identified with the ring generated by $E_{2}$ over the ring of classical holomorphic modular forms $M(\Gamma)$ of integral weight for $\Gamma$ [Zag08, Prop 20(ii)], and by [Zag08, Prop 20(i)] it is stable under the operator
$\theta=q\cdot\frac{d}{dq}=\frac{1}{\pi i}\cdot\frac{d}{d\tau}.$
Let $M^{!}(\Gamma)$ denote the ring of weakly holomorphic modular forms for $\Gamma$; that is, the meromorphic modular forms which are holomorphic away from the cusps.
Definition 7.4.1.
The ring of weakly holomorphic quasi-modular forms $\widetilde{M}^{!}(\Gamma)$ for $\Gamma\subset\mathrm{SL}_{2}(\mathbf{Z})$ is the ring $\widetilde{M}(\Gamma)[1/\Delta]$.
Lemma 7.4.2.
The ring $\widetilde{M}^{!}(\Gamma)$ is the smallest ring which contains $M^{!}(\Gamma)$ and which is closed under $\theta$.
Proof.
If $f\in M^{!}(\Gamma)$ then $f\Delta^{m}\in M(\Gamma)\subset\widetilde{M}(\Gamma)$ for some $m$ and thus $M^{!}(\Gamma)\subset\widetilde{M}^{!}(\Gamma)$. Recall that $\theta\Delta=E_{2}\Delta$. If $g\in\widetilde{M}^{!}(\Gamma)$, then $h=\Delta^{m}g\in\widetilde{M}(\Gamma)$ for some $m$, and thus
$\theta g=\theta\frac{h}{\Delta^{m}}=\frac{\theta h}{\Delta^{m}}-\frac{mE_{2}h}{\Delta^{m}}\in\widetilde{M}^{!}(\Gamma),$
and hence $\widetilde{M}^{!}(\Gamma)$ is closed under $\theta$. Finally, any ring containing $M^{!}(\Gamma)$ and closed under $\theta$ contains both $\Delta^{-1}$ and $E_{2}=\theta\Delta/\Delta$ and thus contains $\widetilde{M}^{!}(\Gamma)$. ∎
Question 7.4.3.
@@
If a component $F_{j}(\tau)$ of a logarithmic vector-valued modular form for $\mathrm{SL}_{2}(\mathbf{Z})$ has a $\mathbf{Z}\llbracket q\rrbracket$ Fourier expansion, does $F_{j}(\tau)$ belong to the ring of weakly holomorphic quasi-modular forms $\widetilde{M}^{!}(\Gamma)$ for some congruence subgroup $\Gamma\subset\mathrm{SL}_{2}(\mathbf{Z})$?
Recall the classical Jacobi theta functions:
$\vartheta_{2}=\sum_{n\in\mathbf{Z}}q^{(n+1/2)^{2}},\quad\vartheta_{3}=\sum_{n\in\mathbf{Z}}q^{n^{2}},\quad\vartheta_{4}=\sum_{n\in\mathbf{Z}}(-1)^{n}q^{n^{2}}.$
By Jacobi’s triple product identity, these functions $\vartheta_{i}$ have explicit representations in terms of the Dedekind $\eta$ function:
$\vartheta_{2}=\frac{2\eta^{2}(2\tau)}{\eta(\tau)},\qquad\vartheta_{3}=\frac{\eta^{5}(\tau)}{\eta^{2}(\tau/2)\eta^{2}(2\tau)},\qquad\vartheta_{4}=\frac{\eta^{2}(\tau/2)}{\eta(\tau)},$
and hence they are holomorphic modular forms of weight $1/2$ without any zeros on $\mathbf{H}$. Consequently, all the Laurent monomials $\vartheta_{2}^{a}\vartheta_{3}^{b}\vartheta_{4}^{c}$ (with $a,b,c\in\mathbf{Z}$) of an even degree $a+b+c$ belong to $\widetilde{M}(\Gamma)$ for some fixed congruence subgroup $\Gamma$ (one can take $\Gamma=\Gamma(12)$, although some monomials are invariant under smaller groups, for example: $(\vartheta_{2}/\vartheta_{3})^{4}=\lambda$ by Equation (1.0.2)). Finally, we also have $2(\theta\vartheta_{i})/\vartheta_{i}=(\theta\vartheta^{2}_{i})/\vartheta^{2}_{i}\in\widetilde{M}^{!}(\Gamma)$.
We now turn to some basic examples hinting towards a positive answer to question 7.4.3. Complementing Example 7.2.3 is the $\lambda$-pullback of the complete elliptic integral of the second kind:
[…]
clearly a component of a logarithmic vector-valued modular form on $\mathrm{SL}_{2}(\mathbf{Z})$, whose $q$-expansion is in $\mathbf{Z}\llbracket q\rrbracket$. But one can indeed verify that
$\frac{2}{\pi}E(\lambda(q))=\frac{\vartheta_{3}\vartheta^{4}_{4}+4\theta\vartheta_{3}}{\vartheta^{3}_{3}}.$
is also an element of $\widetilde{M}^{!}(\Gamma)$ for the congruence subgroup $\Gamma=\Gamma(12)$. One can express $E$ in terms of $K$ and its integral:
$\frac{2}{\pi}E(16x)=(16x-1)\,\frac{2}{\pi}K(16x)-8\int_{0}^{x}\frac{2}{\pi}K(16t)\,dt,$
where one finds that
$\int_{0}^{x}\frac{2}{\pi}K(16t)\,dt=\sum_{n=0}^{\infty}\frac{1}{n+1}\binom{2n}{n}^{2}x^{n+1}\in\mathbf{Z}\llbracket x\rrbracket,$
the integrality of the coefficients now manifested by the Catalan numbers $C_{n}=\frac{1}{n+1}\binom{2n}{n}\in\mathbf{Z}$. One further integration still has integer coefficients:
$\int_{0}^{x}\frac{dy}{y}\int_{0}^{y}\frac{2}{\pi}K(16t)\,dt=\sum_{n=0}^{\infty}\frac{1}{(n+1)^{2}}\binom{2n}{n}^{2}x^{n+1}=\sum_{n=0}^{\infty}C_{n}^{2}\,x^{n+1}\in\mathbf{Z}\llbracket x\rrbracket,$
and $\sum\frac{1}{(n+1)^{2}}\binom{2n}{n}^{2}(\lambda/16)^{n}\in\mathbf{Z}\llbracket q\rrbracket$ is a component of a logarithmic vector-valued modular form with a $\mathbf{Z}\llbracket q\rrbracket$ expansion. Zudilin has pointed out to us the formula
$\sum_{n=0}^{\infty}\frac{1}{(n+1)^{2}}\binom{2n}{n}^{2}\big(\lambda(q)\big/16\big)^{n}=\frac{4}{\vartheta^{4}_{2}}\left(4\vartheta^{2}_{3}\cdot\frac{\theta\vartheta_{2}}{\vartheta_{2}}+4\vartheta_{3}\cdot\theta\vartheta_{3}-\vartheta^{4}_{3}\right)$
exhibiting this $\mathbf{Z}\llbracket q\rrbracket$ power series as an element of $\widetilde{M}^{!}(\Gamma)$ for the congruence subgroup $\Gamma=\Gamma(12)$, in accordance with Question 7.4.3.
@@
8. Acknowledgments
We would like to thank Yves André for a number of insightful remarks on the first version of this manuscript, leading in particular to § 2.4.1 as an alternative and simplified approach to Theorem 2.0.1. We would also like to thank Jean-Benoît Bost, Pierre Deligne, Cameron Franc, Javier Fresán, Mark Kisin, Geoffrey Mason, Alex Smith, Richard Taylor, John Voight, and Wadim Zudilin for useful remarks, suggestions, and corrections.